OpenStudy (mwhit2101):

SIXTH GRADE MATH HELP! MEDAL + FAN!! Write an equation in slope-intercept form for the line passing through the pair of points. (-1,6) (3, -6)

OpenStudy (mwhit2101):

So I know we have to get slope first and the slope =6. So what now.

OpenStudy (eric_d):

use this formula \[y-y _{1}=m(x-x _{_{1}})\]

OpenStudy (mwhit2101):

I'm confused with which one is y1 and y2 and which one is x1 and x2?

OpenStudy (mwhit2101):

x1=-1 y1=6 x2=3 y2=-6. Is this ^^^^ right?

rishavraj (rishavraj):

u r wrong... its slope isn't 6

OpenStudy (eric_d):

u can choose any of them (-1,6) or the other one

OpenStudy (mwhit2101):

-4? um ok

OpenStudy (eric_d):

m=y2-y1/x2-x1

OpenStudy (mwhit2101):

Wait in your forumala what is just plain "y" is that supposed to be y2

OpenStudy (eric_d):

substitute the value

OpenStudy (eric_d):

that's the second step

OpenStudy (eric_d):

to find the final answer

OpenStudy (mwhit2101):

omg im so confused

OpenStudy (mwhit2101):

and also i just did the slope based on yoru formula thing and its -3

OpenStudy (mwhit2101):

x1=-1 y1=6 x2=3 y2=-6 y2-y1/x2-x1= -6-1 3-(-1) -12/4=====-3.

OpenStudy (eric_d):

okay sorry it's -3 overlook

OpenStudy (mwhit2101):

So how do I get an equation

OpenStudy (eric_d):

continue with the other formula

OpenStudy (mwhit2101):

yah I plugged everything in and I got: -6-6=-3(3-(-1)) and its true according to my calculator

OpenStudy (eric_d):

this is definitely wrong

OpenStudy (eric_d):

choose only the first pair

OpenStudy (mwhit2101):

Ok do you know the answer or not because I'm really confused right now and we don't seem to be getting anywhere and I'm in the middle of my math final exam and I really need an answer like now.

OpenStudy (eric_d):

ok

OpenStudy (eric_d):

-1,6 y-6=-3(x+1) y=1/2x+1/2

OpenStudy (mwhit2101):

Those are non of my choices

OpenStudy (mwhit2101):

A. y=-3x-3 B y=3x+3 C y=-3x+3

OpenStudy (mwhit2101):

D. y=1/3x-3

OpenStudy (eric_d):

B & D r incorrect coz m not =-3

OpenStudy (mwhit2101):

Yes...

OpenStudy (mwhit2101):

Oh I know!

OpenStudy (mwhit2101):

It's A probably because the slope is -3 remember?

OpenStudy (eric_d):

c is also -3

OpenStudy (mwhit2101):

no, C is +3

OpenStudy (eric_d):

@ganeshie8